### v4

# Low-Confidence Gold: Refining Low-Confidence Samples for Efficient Instruction Tuning

Abstract
The effectiveness of instruction fine-tuning for Large Language Models is fundamentally constrained by the quality and efficiency of training datasets. This work introduces Low-Confidence Gold (LCG), a novel filtering framework that employs centroid-based clustering and confidence-guided selection for identifying valuable instruction pairs. Through a semi-supervised approach using a lightweight classifier trained on representative samples, LCG curates high-quality subsets while preserving data diversity. Experimental evaluation demonstrates that models fine-tuned on LCG-filtered subsets of 6K samples achieve superior performance compared to existing methods, with substantial improvements on MT-bench and consistent gains across comprehensive evaluation metrics. The framework’s efficacy while maintaining model performance establishes a promising direction for efficient instruction tuning.
Low-Confidence Gold: Refining Low-Confidence Samples for Efficient Instruction Tuning
Hongyi Cai1, Jie Li2, Mohammad Mahdinur Rahman1, Wenzhen Dong3 1Universiti Malaya, 2University of Science and Technology Beijing, 3The Chinese University of Hong Kong Correspondence: xcloudfance@gmail.com
1 Introduction
Large Language Models (LLMs) have been trained to follow instructions by specific supervised response data after pre-training stage. Many instruction finetuning (IFT) Taori et al. (2023) datasets emerge to realize various downstream tasks, for example: mathematic calculation, sentence analysis, haiku writing and etc, aiming to strengthen the ability of LLMs in instruction following. To save vast human costs for data annotation, most of studies introduce other teacher LLMs (e.g. text-davinci-003 Brown et al. (2020)) to align the best instructions with corresponding responses.
However, IFT datasets (e.g. Alpaca_52k Taori et al. (2023), magpie Xu et al. (2024)) suffer from misleading content and poor quality, resulting in the bottleneck of post-training performance, even though teacher models replenish the missing parts of context and instruction pairs. This highlights the need for effective data filtering methods that identify high-quality instruction subsets while reducing fine-tuning time and computational costs.
Figure 1: We target to select complex and quality samples confidence ranking for benefiting LLM training.
Alpagasus Chen et al. (2024) proposed a model-based approach that introduces proprietary LLMs to score data quality in multiple facets, replacing human annotation by taking advantage of the automated pipeline. However, this leads to datasets that are likely biased by the preference for redundant and limited responses Panickssery et al. (2024), which potentially deteriorates the diversity of the original data. Ge et al., 2024 emphasizes the necessity of diversity and therefore proposed clustering and ranking to select subsets of data. Further, Superfiltering Li et al. (2024) gains more insights in small open-source LLM that scores the instruction following ability of Alpaca_52k. Although the instruction score provides an efficient and simple criterion for data selection, it does not consistently correlate with both the quality and diversity of data. Consequently, improvements in performance may not always be guaranteed.
To address these challenges, we propose a novel data filtering framework, Low-Confidence Gold (LCG) for efficient instruction tuning that significantly reduces computational costs while maintaining model performance. Our approach, shown in 1, innovatively seeks to identify high-value instruction data through classification tasks. Specifically, we develop a lightweight classification model trained on centroid subsets that effectively categorizes instruction-response pairs, and leverage low-confidence predictions to curate challenging examples most beneficial for instruction tuning. Another perspective is that, since the common instruction tuning data are lack of annotations and labels, we adopt the manner of semi-supervised learning, to construct pseudo-labels as our training groundtruth, as well as getting inspired quality data from affordable yet effective models.
Through extensive experiments on the Alpaca_52K dataset, we demonstrate that our filtered subsets achieve comparable or better performance when fine-tuning various open-source language models, while requiring only a fraction of the original data. Our main contributions are threefold:
  1. 1.
    A novel and efficient data filtering paradigm for instruction tuning that combines nearest neighbor classification with confidence-based selection.
  2. 2.
    We train a small classifier model that enables selection for the whole set of instruction finetuning data.
  3. 3.
    Experiments and evaluations are conducted that demonstrate the outstanding effectiveness of our filtered datasets working on multiple open-source LLMs. We reach states-of-the-arts performance in MT-Bench and HuggingFace OpenLLM Leaderboard benchmarks.
2 Preliminaries
2.1 K-means Clustering
Given the Alpaca_52k dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ where $N=52,000$, we first cluster instructions into $K$ semantic groups using K-means. Let $\phi(x_{i})\in\mathbb{R}^{d}$ denote the embedding vector of instruction $x_{i}$. The clustering objective minimizes:
$\min_{\{C_{k}\}_{k=1}^{K}}\sum_{k=1}^{K}\sum_{x_{i}\in C_{k}}\|\phi(x_{i})-\mu_{k}\|^{2}$ (1)
where $\mu_{k}=\frac{1}{|C_{k}|}\sum_{x_{i}\in C_{k}}\phi(x_{i})$ is the centroid of cluster $C_{k}$. This partitions $\mathcal{D}$ into $K$ disjoint subsets $\{C_{1},...,C_{K}\}$ based on instruction similarity.
2.2 Problem Setting
Our filtering framework, LCG aims to select a subset $\mathcal{D}_{\text{filtered}}\subseteq\mathcal{D}$ that satisfies:
$\mathcal{D}_{\text{filtered}}=\bigcup_{k=1}^{K}\left\{(x_{j},y_{j})\in C_{k}\mid\mathcal{F}(x_{j},y_{j})<\tau_{k}\right\}$ (2)
where $\mathcal{F}:\mathcal{D}\rightarrow[0,1]$ is a discriminative confidence scorer and $\tau_{k}$ is an adaptive threshold for cluster $C_{k}$. The scorer $\mathcal{F}$ evaluates how "hard" a sample is to be trivially categorized, with higher values indicating the simplicity of data which is easily determined and differentiated. The training efficiency therefore increases since only a small subset of instructions are curated.
Figure 2: The overall pipeline of Low-Confidence Gold. We split our pipeline into two main steps: 1) Clustering to get pseudo-labels and centroid data to collect the initial diversity of data. 2) We feed annotated data into a tiny yet effective classifier to rank the confidences for the rest of the distant data to implement subset selection.
3 Methodology
3.1 Motivation
Instruction filtering demands a dual-focus mechanism that intrinsically balances data quality and diversity. Traditional supervised methods face inherent scalability limitations as manual annotation becomes prohibitively expensive for large-scale instruction datasets Liu et al. (2022); Longpre et al. (2023); Liu et al. (2023). Meanwhile, it is difficult to identify suitable and challenging data for LLMs training without introducing proprietary LLMs or labors. Our semi-supervised framework addresses these limitations through pseudo-label refinement and early-stopped confidence detection, creating dynamic selection boundaries aligned with language model learning dynamics.
Cluster-centric pseudo-labeling addresses data distribution challenges in instruction tuning. Traditional sampling methods often struggle to balance between common and rare instruction patterns, leading to either over-representation of frequent cases or loss of valuable rare examples. We create semantic clustering anchors that naturally preserve the diversity of instruction patterns. By sampling 3% of data points nearest to cluster centroids, we ensure each semantic category contributes meaningful examples while maintaining the inherent data distribution characteristics.
Early-stopped classifier training induces uncertainty to identify high-quality samples. Limiting the classifier to 3 epochs creates deliberate underfitting - the model develops basic pattern recognition without over-specializing to pseudo-labels. When applied to non-centroid samples, this partially-trained classifier’s low-confidence predictions signal instructions containing non-trivial semantic constructs. These samples challenge the classifier’s emerging decision boundaries precisely because they contain valuable complexity that language models should master, not avoid.
Model MT-bench Huggingface Open LLM Leaderboard Scores (%)
Score Hellaswag MMLU GSM8k ARC Avg
First Group - Base Model
Mistral-7b-v0.3 3.639 60.94 58.96 36.62 48.81 51.33
First Group - Methods
Alpaca-52k 4.018 61.18 57.73 31.61 53.07 50.90
SuperFiltering-10% 3.963 60.98 59.34 35.71 49.83 51.47
Random-6k 4.314 60.83 58.75 35.03 53.07 51.92
Perplexity-6k 4.352 61.64 58.48 37.00 51.88 52.25
Kmeans-6k 4.283 60.86 58.45 35.10 52.05 51.62
LIMA-6k 4.440 60.58 59.34 37.31 51.11 52.09
LCG-MultinomialNB-6k (Ours) 5.086 62.00 59.51 40.51 52.90 53.73
LCG-DistilBERT-6k (Ours) 4.894 61.99 59.51 40.33 52.22 53.51
LCG-DistilBERT-1k (Ours) 4.869 61.94 59.24 38.29 51.62 52.77
Second Group - Base Model
LLaMa3-8b 3.418 60.17 62.13 50.42 50.26 49.98
Second Group - Methods
Alpaca-52k 3.718 60.57 61.36 46.10 52.41 55.74
SuperFiltering-10% 3.968 60.38 61.95 50.34 51.54 55.36
Random-6k 3.912 60.83 58.75 35.03 53.07 51.92
Perplexity-6k 4.120 61.14 61.09 50.87 53.50 56.65
Kmeans-6k 3.731 60.86 58.45 35.10 53.07 51.87
LIMA-6k 4.450 60.58 62.13 50.34 51.11 55.82
LCG-DistilBERT-6k (Ours) 4.963 61.43 62.67 54.28 54.78 58.29
LCG-DistilBERT-1k (Ours) 4.776 60.95 62.26 52.92 52.82 57.23
Table 1: Performance comparison on standard benchmarks. Results in bold indicate best performance within each group, while underlined values represent second-best performance within each group. The table is divided into two groups, each with its base model and various fine-tuning methods.
3.2 Centroid Coreset Selection for Pseudo-labels
In the initial step of our approach, we select a coreset from the whole corpus to identify pseudo-labels by the K-means algorithm, which effectively determine each semantic clusters. Given a dataset of instruction pairs $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, we first encode each instruction $\mathbf{x}_{i}$ into a dense vector representation using MiniLM Wang et al. (2022):
$\mathbf{h}_{i}=\text{AvgPool}(\text{MiniLM}(\mathbf{x}_{i}))\in\mathbb{R}^{384}$ (3)
This geometric progression ensures proportional coverage of both frequent and rare instruction patterns. Cluster centroids $\{\mathbf{c}_{j}\}_{j=1}^{k}$ are computed via:
$\mathbf{c}_{j}=\frac{1}{|\mathcal{C}_{j}|}\sum_{\mathbf{x}_{i}\in\mathcal{C}_{j}}\tilde{\mathbf{h}}_{i}$ (4)
where $\mathcal{C}_{j}$ denotes the set of samples assigned to cluster $j$. Centroid-proximal samples are selected as high-confidence candidates:
$\mathcal{D}_{\text{core}}=\{\mathbf{x}_{i}|\|\tilde{\mathbf{h}}_{i}-\mathbf{c}_{j(i)}\|_{2}<\gamma\}$ (5)
where $\gamma$ is the 90th percentile distance within each cluster.
3.3 Low-Confidence Gold: Calibrating with Low-confidence samples to select data
After determining pseudo-labels based on clusters, those annotations can be served for classification training. Specifically, we train a multi-class classifier on the core samples $\mathcal{D}_{\text{core}}$. The model architecture consists of:
$f_{\theta}(\mathbf{x})=\text{Softmax}(\mathbf{W}_{2}\cdot\text{GELU}(\mathbf{W}_{1}\mathbf{h}_{i}+\mathbf{b}_{1})+\mathbf{b}_{2})$ (6)
where $\mathbf{W}_{1}\in\mathbb{R}^{384\times 768}$, $\mathbf{W}_{2}\in\mathbb{R}^{768}$ are learnable parameters, and GELU denotes the Gaussian Error Linear Unit activation. The model optimizes cross-entropy loss:
$\mathcal{L}(\theta)=-\frac{1}{|\mathcal{D}_{\text{core}}|}\sum_{(\mathbf{x}_{i},y_{i})}$ $\sum_{j=1}^{k}\mathbb{I}(y_{i}=j)$ (7)
$\cdot\log p_{\theta}(y=j|\mathbf{x}_{i})$
Training terminates at epoch $T=3$ since we aim to keep the model in an early-stopped stage so that they would not overfit to the centroid subset data. After training, we rank the confidence distribution calculated from softmax function and select the top $K$ most uncertain data in each cluster.
4 Experiments
In this section, we utilize LCG to filter Alpaca_52k dataset into 6k and evaluate the subset by fine-tuning in 2 open-source LLMs: 1) Mistral-7b-v0.3 Jiang et al. (2023) and 2) LLaMa3-8b Dubey et al. (2024). Additionally, we adopt two different classifiers as our semi-supervised training model to examine difficult samples. Both results demonstrate the effectiveness of our method, as shown in Tab. 1.
Settings. Two classifiers are Multinomial Naive Bayes Pedregosa et al. (2011) and DistilBERT Sanh et al. (2019) respectively. We set the training learning rate as 1e-5 and also 3 epochs to train as mentioned before. After training is finished, we select curated datasets by confidences < 0.7 to fine-tune open-source LLMs by LoRA Hu et al. (2022) with a learning rate of 2e-5 and 3 epochs.
Our proposed Low-confidence Gold (LCG) method consistently outperforms existing instruction data filtering approaches across multiple base models and evaluation benchmarks. When applied to Mistral-7b, LCG with MultinomialNB achieves the highest MT-bench score of 5.086, surpassing the previous best (LIMA-6k Zhao et al. (2024)) by 14.5%. Similarly, LCG with DistilBERT demonstrates superior performance on LLaMA3-8b, improving the MT-bench score by 11.5% over LIMA-6k. Notably, our method maintains strong performance even with only 1k examples, highlighting its effectiveness in identifying high-quality instruction data. The consistent improvements across diverse metrics (Hellaswag Zellers et al. (2019), MMLU Hendrycks et al. (2021), GSM8k Cobbe et al. (2021), and ARC Clark et al. (2018)) further validate the robustness of our approach.
5 Conclusion
In this paper, we proposed Low-Confidence Gold (LCG), a novel data filtering framework that combines cluster-centric pseudo-labeling with early-stopped classifier training for efficient instruction tuning. Through extensive experiments, we demonstrated the strong performance across multiple benchmarks and base models, validating the effectiveness of our semi-supervised learning paradigm in maintaining both data quality and diversity for instruction tuning.
6 Limitation
Our work introduces a semi-supervised training paradigm to curate a subset of data for instruction tuning based on confidence score. However, there still exist several challenges: 1) Even though classifiers are tiny and spend low computational resources to train, it still takes time and effort to initially select data with annotated pseudo-labels. 2) It is likely to be hindered by the original biases and tasks of the dataset, which might still cause inefficiency after selection.
Appendix A Extended Analysis of Semi-Supervised Model Configurations
A.1 MultinomialNB Implementation
The confidence distribution patterns of our MultinomialNB baseline, as visualized in Fig. 4, reveal fundamentally different characteristics compared to deep learning architectures. The histogram demonstrates remarkable uniformity across confidence intervals (0.0-1.0 with 0.1 increments), showing no significant concentration in specific confidence ranges. This equilibrium phenomenon stems from the model’s inherent probabilistic nature and linear decision boundaries, which produce well-calibrated confidence estimates despite its simplicity.
A.2 DistilBERT comparative experiment on learning rate
Our DistilBERT implementation employed a systematic exploration of learning rate hyperparameters 1e-4, 1e-5, 1e-6 within the following experimental framework:
  1. 1.
    Architecture: DistilBERT-base-uncased (66M parameters) with custom classification head.
  2. 2.
    Optimization: Adam optimizer.
  3. 3.
    Training regime: 3-epoch constraint to prevent overfitting in low-data scenarios.
  4. 4.
    Data alignment: Identical train/test splits (stratified sampling) as MultinomialNB for direct comparability.
The empirical results (shown in Fig. 4) demonstrate non-monotonic performance relationships with learning rate scaling. Peak accuracy (62%) emerged at 1e-5, while extreme values at both ends (1e-4: 36%, 1e-6: 28%) showed substantial performance degradation. This U-shaped accuracy curve suggests the existence of optimal learning rate basins in semi-supervised BERT fine-tuning.
The model exhibited distinct confidence distribution characteristics at the 1e-6 learning rate, with predictions predominantly clustered in the low-confidence range (0-0.2). However, as revealed in Figure 2, comparative analysis across learning rates demonstrated minimal performance variation, showing only marginal improvements that correlated with accuracy increments.
Figure 3: The data distribution of MultinomialNB across different confidence intervals.
Figure 4: The data distribution of DistilBERT across different confidence intervals under various learning rates.